OpenStudy (anonymous):

what is the derivative of this f(x)=2x−5/6x−8.

OpenStudy (anonymous):

\[\frac{2x-5}{6x-8}\]

OpenStudy (anonymous):

quotient rule for this one

OpenStudy (anonymous):

\[f'(x)=\frac{6x-8)\times 2-(2x-5)\times 6}{(6x-8)^2}\] then some algebra to clean up the numerator

OpenStudy (anonymous):

\[\frac{(2x-5)\prime \times(6x-8)-(2x-5)(6x-8)\prime}{(6x-8)^{2}}=\frac{2(6x-8)-6(2x-5)}{(6x-8)^{2}}\]

OpenStudy (anonymous):

thanks